OpenStudy (fellowroot):

how do i find an inverse trig angle that is not on the unit circle and without using a calculator. For example: arcsec(3/5) = ?

OpenStudy (campbell_st):

well this is a really bad example but here is the logic \[\sec = \frac{1}{\cos}\] so the question is really saying \[\sec(x) = \frac{3}{5} \] or \[\frac{1}{\cos(x) }= \frac{3}{5} \] take the reciprocal of both sides \[\cos(x) = \frac{5}{3}\] now you could find the angle by unfortunately it doesn't exist... since the cos of an angle is always less than 1. so before finding x = arcsec(5/3) use some basics. sec(x) = 5/3 cos(x) = 3/5 x = arccos(3/5) this would work.

OpenStudy (fellowroot):

Thanks, but then how would you find x = arccos(3/5) with no calculator and no unit circle

OpenStudy (campbell_st):

well one method would be approximate \[\cos(45) = 1/\sqrt{2} = 0.707\] cos(60) = 1/2 so arccos(0.6) is about half way between the 2 angles 45 and 60... so approximate it at 52.5 degrees. or maybe 53 use the exact values to help approximate

OpenStudy (campbell_st):

you also know the values for sin, cos and tan of 0 degrees, 30 degrees and 90 degrees... the will assist in approximating. Thats my best idea... other than looking at a trig table.

OpenStudy (fellowroot):

Thank you campbell, but I still need to know without having to approximate.

OpenStudy (phi):

In general, you need a calculator. There are only some small subset of angles that you can figure out using other methods.

OpenStudy (mertsj):

first of all, the secant of an angle cannot be 3/5 because the secant is always greater than 1 or less than negative 1 so the answer to that question is does not exist.

OpenStudy (mertsj):

Secondly, if you had the secant of an angle that did exist, let us say, perhaps, 5/3, you could draw the triangle and find all 6 functions of that angle but if you wanted the actual angle, you would need either a table of values or a calculator.

OpenStudy (fellowroot):

So you are telling me that for some sides you simply cannot find an actual angle without a calculator?

OpenStudy (mertsj):

yes.

OpenStudy (mertsj):

That's why, in the old days, before calculators, we used to carry books full of tables around with us.